# Measuring multipartite quantum correlations by thermodynamic work extraction

### v1

@@ III.2 Upper and lower bounds for work deficit density
where $I_{l}\equiv\{ml|m\in\mathbbm{Z},0\leq m<\frac{N}{l}\}$ is the set of the party labels. Here, $\mathcal{M}_{n:l}$ represents the set of measurement protocols composed of the local projective measurements on $\{a_{i}\}_{i=n+1}^{n+l}$, and $Y_{n:l}\equiv(y_{n+1},\dots,y_{n+l})$ represents their outcomes, as shown in Fig. 6(b)  [^1] We note that for $n$ such that $N-l+1\leq n<N$, the definitions of $\mathcal{M}_{n:l}$ and $Y_{n:l}$ are changed reflecting the existence of the edge in the system: $\mathcal{M}_{n:l}$ is the set of protocols where the local measurements on $\{a_{i}\}_{i=n+1}^{N}$ is performed, and $Y_{n:l}\equiv(y_{n+1},\dots,y_{N})$ denotes the corresponding outcomes.. By denoting the measurement outcomes from parties $\{a_{i}\}_{i=1}^{n}$ as $Y_{n}\equiv(y_{1},\dots,y_{n})$ this upper bound can be derived as follows:
@@
In the upper bound $U_{l}$ defined as Eq. (16), the global optimization of $\delta$ is divided into the optimization of $l$-site segments $\min_{\Lambda}H_{\Lambda}(Y_{n:l})$. The cost for optimizing each term $\min_{\Lambda}H_{\Lambda}(Y_{n:l})$ is $e^{\mathcal{O}(l)}$, due to the necessity to optimize the local measurement bases on $l$ sites. Since $U_{l}$ is calculated as the sum of $\frac{N}{l}$ such terms, the total computational cost is represented as $Ne^{\mathcal{O}(l)}$, capturing the dominant factors of $N$ and $l$. Furthermore, for the translationally invariant infinite system described with the normal MPS, $U_{l}$ can be calculated with a cost of $e^{\mathcal{O}(l)}$. This is because almost all the terms $\min_{\Lambda}H_{\Lambda}(Y_{n:l})$ in $U_{l}$ are reduced to the same value due to the translational invariance and finite correlation length, eliminating the need for summation. More detailed discussion is provided in Supplemental Material.
@@
where the inequality $\delta\leq U_{\Lambda^{*}}$ directly follows from the definition of $\delta$. In the calculation for the AKLT state and the cluster state in Sec. III.3.1 and III.3.2, we employ this type of upper bound by choosing appropriate ansatz protocols $\Lambda^{*}$. As another type of upper bound, we introduce $U_{l,k}$ with natural numbers $l$ and $k$, which is derived by restricting the optimization region of the measurement protocols. This bound $U_{l,k}$ is the generalization of the upper bound $U_{l}$, and is utilized in the numerical calculation in Sec. III.3.3. The definition and derivation of this type of bound are provided in Appendix D.
@@
where $\mathcal{N}_{n:l}$ denotes the set of measurement protocols composed of arbitrary POVM on the sites $\{a_{i}\}_{i=1}^{n}$ and local projective measurements on sites $\{a_{i}\}_{i=n+1}^{n+l}$  [^2] For $n$ such that $N-l+1\leq n<N$, $\mathcal{N}_{n:l}$ becomes the set of protocols where the POVM on $\{a_{i}\}_{i=1}^{n}$ and the local measurements on $\{a_{i}\}_{i=n+1}^{N}$ are performed., as shown in Fig. 6(c). Here, $Z_{n}$ denotes the outcome of this POVM. By denoting the set of the local projective measurement on $\{a_{i}\}_{i=1}^{n+l}$ as $\mathcal{M}_{n+l}$, this bound is derived as
@@
From Eq. (19), we can see that this lower bound $L_{l}$ is defined as the sum of the local terms $\min_{\Lambda}H_{\Lambda}(Y_{n:l}|Z_{n})$, which is similar to the upper bound $U_{l}$. The difference is that in the lower bound $L_{l}$, we need to optimize the POVM on the $d^{n}$ dimensional subsystem $\{a_{i}\}_{i=1}^{n}$, whose numerical cost seems to grow exponentially with the subsystem size $n$. However, in the MPS with fixed bond dimension $D_{B}$, we can perform this optimization with the constant numerical cost, since the subsystem $\{a_{i}\}_{i=1}^{n}$ can be effectively regarded as $D_{B}$-dimensional system. Because of this reduction of subsystem dimension, the numerical cost for $L_{l}$ becomes $Ne^{\mathcal{O}(l)}$ up to constant multiple in $N$-site MPS, and $e^{\mathcal{O}(l)}$ for the translationally invariant infinite systems described with normal MPS.
@@
where $k$ is arbitrary natural number. Furthermore, for systems described with MPS, we can derive Eq. (14), which indicates that an arbitrary error bound is achieved with our method by increasing $l$. Since the numerical costs for calculating the bounds $L_{l}$ and $U_{l}$ scales as $e^{\mathcal{O}(l)}$ with the locality parameter $l$, Eq. (14) demonstrates that the cost scales with the inverse of the error bound, $\varepsilon$, as $e^{\mathcal{O}({\frac{1}{\varepsilon}})}$ (Table 1).
@@ III.3 Demonstration in examples
We demonstrate our efficient calculation method introduced in Sec. III.1 and III.2 in some examples. Especially, we calculate the work deficit density in translationally invariant infinite MPS, where the direct calculation is impossible, while our calculation method can be performed with reasonable numerical cost.
@@ III.3.1 AKLT state
We first calculate the work deficit density for the AKLT state Affleck et al. (1987, 1988) in infinite one-dimensional system under open boundary condition. The AKLT state is the ground state of the AKLT Hamiltonian
@@
$A^{\uparrow}=\sqrt{\frac{2}{3}}\begin{pmatrix}0&0\\ -1&0\end{pmatrix},A^{0}=\sqrt{\frac{1}{3}}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},A^{\downarrow}=\sqrt{\frac{2}{3}}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$ (22)
While the direct calculations of $\delta$ is infeasible since we need to optimze infinite sets of bases, the calculation of the lower bound $L_{1}$ can be performed with the optimization in $2\times 3$ dimensional system, due to the fixed bond dimension $D_{B}=2$. In this case, we can evaluate the lower bound through analytical calculation, as
@@
Therefore, we can say that the multipartite work deficit in the AKLT state becomes the relative entropy of coherence with respect to the orthogonal basis $\{\otimes_{n}\ket{i_{n}}_{n}\}_{i_{n}}$, where |$\ket{i_{n}}_{n}$ denotes the basis of $a_{n}$ chosen from $i_{n}=\uparrow,0,\downarrow$.
@@ III.3.2 Cluster state
We next calculate the work deficit density in the cluster state Briegel and Raussendorf (2001) in one-dimensional infinite system under open boundary condition. The cluster state is the ground state of the cluster Hamiltonian
@@
$A^{0}=\sqrt{\frac{1}{2}}\begin{pmatrix}0&0\\ 1&1\end{pmatrix},A^{1}=\sqrt{\frac{1}{2}}\begin{pmatrix}1&-1\\ 0&0\end{pmatrix}.$ (26)
While the direct calculation is not applicable for infinite systems, we can calculate the lower bounds $L_{2}$ by performing the optimization over $2\times 2\times 2$ dimensional system. This optimization can be performed analytically in this case, and the lower bound is obtained as
@@
Since the lower bound of a single site segment is calculated as $L_{1}=0$ in this state, only providing the trivial lower bound, we can see that the bound is significantly improved by increasing locality parameter $l$.
We can also derive the upper bound $U_{\Lambda^{*}}$ with the ansatz protocol $\Lambda^{*}$ where the projective measurement on {|$\{\ket{0},\ket{1}\}$ and {$\{\ket{+},\ket{-}\}$ are alternatively performed in each site. Here, basis $\ket{\pm}$ are defined as |±$\ket{\pm}\equiv\frac{1}{\sqrt{2}}(\ket{0}\pm\ket{1})$. This upper bound is calculated as
@@
From these calculations, we can see that the multipartite work deficit for the cluster state becomes the relative entropy of coherence with respect to the basis $\{\otimes_{n}\ket{i_{n}}_{2n-1}\ket{j_{n}}_{2n}\}_{i_{n},j_{n}}$, where $i_{n}$ is chosen from $i_{n}=0,1$ and $j_{n}$ is chosen from $j_{n}=\pm$.
@@ III.3.3 MPS family under quantum phase transition
Figure 7: Correlation length as a function of the system parameter $g$ for an MPS family (30). The correlation length diverges at the quantum phase transition point $g=0$, indicating a transition from a symmetry-protected topological (SPT) phase when $g<0$ to a trivial phase when $g>0$.
@@
We consider the region $-1\leq g\leq 1$, and infinite system under preodic boundary condition. These MPS are the ground states of the parametrized Hamiltonian
@@
From Fig. 8(a), we can observe that the bimodal structure of $\delta$, with its peaks in the regions $-1\leq g\leq 0$ and $0\leq g\leq 1$, respectively. The reasons for this bimodal structure can be attributed to two factors: long-range correlation around $g=0$ and the short-range correlation around $g=\pm 1$. At $g=1$, work deficit density $\delta$ becomes zero, since the system is in a product state and does not have inter-site correlation at all. On the other hand, at $g=0$, $\delta$ drop to zero due to the non-decaying correlation. The generalized GHZ state realized at $g=0$, is the superposition of the completely aligned states |$\ket{00...0}$ and |$\ket{11...1}$, and therefore the multipartite work deficit remains the same value $\Delta=\ln 2$ independent of the system size $N$. This leads to $\delta\equiv\frac{\Delta}{N}=0$ at $g=0$ in the infinite system $N\to\infty$. Consequently, the peak of $\delta$ occurs between these two extremes, where the system has the intermediate-scale correlation length. The peak in the region $-1\leq g\leq 0$ occurs in a similar mechanism.
We further discuss the kinks observed in Fig. 8(a). In this figure, the unsmooth change in $\delta$ is observed at the phase transition point $g=0$, reflecting the drastic change of the quantum correlation structure. Around this point $g=0$, the optimal measurement protocol $\Lambda^{*}$ is kept unchanged; it is to perform the measurement on the basis {|$\{\ket{0},\ket{1}\}$ for all sites. On the other hand, the kink observed around $g=0.175$ in Fig. 8(a) is due to a discontinuous change in the optimal measurement protocol at this point; in the region $g<0.175$, the optimal protocol is to perform projective measurements on the basis {|$\{\ket{0},\ket{1}\}$ for all sites, while for $g\geq 0.175$, the optimal basis suddenly switches to {$\{\ket{+},\ket{-}\}$, where |±$\ket{\pm}\equiv\frac{1}{\sqrt{2}}(\ket{0}\pm\ket{1})$. Additionally, the kink may occur for the same reason in the region $\-1\leq g\leq 0$, although it is not clearly observable in the current numerical calculation due to the gap in the lower and upper bounds. Such kinks, resulting from changes in the optimal measurement protocols, do not necessarily indicate a drastic change in the quantum correlation structure at those points. To distinguish between such kinks and the unsmoothness due to the phase transition at $g=0$, we further introduce the framework of coarse-graining in Sec. IV, demonstrating that the kink around $g=0$ becomes sharper while the kinks at the other points get dull under the caorse-graining.
@@ IV Coarse-graining
Figure 9: Schematic for the coarse-graining. In this figure, the neighboring three sites $\{a_{3n-2},a_{3n-1},a_{3n}\}$ are merged and regarded as a single party $b_{n}$.
In the foregoing sections, we calculated the work deficit density $\delta$ defined in the partitioning where each site is regarded as a single party. In this section, we discuss $\delta$ defined in the coarse-grained grouping. To clarify the partitioning under which $\delta$ is defined, we use the following notation in this section:
$\delta_{\{a_{n}\}_{n=1}^{N}},$
where the subscript $\{a_{n}\}_{n=1}^{N}$ denotes the grouping, while it can be abbreviated in the obvious cases.
In the framework of coarse-graining, we calculate the work deficit density defined under the grouping that several parties are merged, as illustrated in Fig. 9. By utilizing this framework, we can get access to the structure of longer-range multipartite quantum correlation. For example, when the neighboring sites $\{a_{3n-2},a_{3n-1},a_{3n}\}$ are merged and regarded as a single party $b_{n}$, the work deficit density defined in this partitioning, $\delta_{\{b_{n}\}_{n=1}^{N/3}}$, does not account for the short-range quantum correlation among parties $\{a_{3n-2},a_{3n-1},a_{3n}\}$, while it reflects the structure of longer-range quantum correlation. This is contrary to $\delta_{\{a_{n}\}_{n=1}^{N}}$, which is defined under the original partitioning and accounts for the correlation among the neighboring sites $\{a_{3n-2},a_{3n-1},a_{3n}\}$. In this section, we discuss the behavior of the work deficit density defined in such a coarse-grained picture, probing the long-range structure of multipartite quantum correlation.

### v2

@@ III.2 Upper and lower bounds for work deficit density
where $I_{l}\equiv\{ml|m\in\mathbbm{Z},0\leq m<\frac{N}{l}\}$ is the set of the party labels. Here, $\mathcal{M}_{n:l}$ represents the set of measurement protocols composed of the local projective measurements on $\{a_{i}\}_{i=n+1}^{n+l}$, and $Y_{n:l}\equiv(y_{n+1},\dots,y_{n+l})$ represents their outcomes, as shown in Fig. 6(b)  [^2] We note that for $n$ such that $N-l+1\leq n<N$, the definitions of $\mathcal{M}_{n:l}$ and $Y_{n:l}$ are changed reflecting the existence of the edge in the system: $\mathcal{M}_{n:l}$ is the set of protocols where local projective measurements on $\{a_{i}\}_{i=n+1}^{N}$ is performed, and $Y_{n:l}\equiv(y_{n+1},\dots,y_{N})$ is the corresponding outcomes.. By denoting the measurement outcomes from parties $\{a_{i}\}_{i=1}^{n}$ as $Y_{n}\equiv(y_{1},\dots,y_{n})$, this upper bound can be derived as follows:
@@
In the upper bound $U_{l}$ defined as Eq. (16), the global optimization of $\delta$ is divided into the optimization of $l$-site segments. The cost for optimizing each term $\min_{\Lambda}H_{\Lambda}(Y_{n:l})$ is $e^{\mathcal{O}(l)}$, due to the necessity to optimize the local measurement bases on $l$ sites. Since $U_{l}$ is calculated as the sum of $\frac{N}{l}$ such terms, the total computational cost is represented as $Ne^{\mathcal{O}(l)}$, capturing the dominant factors of $N$ and $l$. Furthermore, for translationally invariant normal MPS in the thermodynamic limit, $U_{l}$ can be calculated with a cost of $e^{\mathcal{O}(l)}$. This is because almost all the terms in the summation $\{\min_{\Lambda}H_{\Lambda}(Y_{n:l})\}_{n\in I_{l}}$ are reduced to the same value due to the translation invariance and finite correlation length, eliminating the need for summation. More detailed discussion is provided as Theorem 3 in Supplemental Material.
@@
where the inequality $\delta\leq U_{\Lambda^{*}}$ directly follows from the definition of $\delta$. In the calculation for the AKLT state and the cluster state in Secs. III.3.1 and III.3.2, we employ this type of upper bound by choosing appropriate ansatz protocols $\Lambda^{*}$. As another type of upper bound, we introduce $U_{l,k}$ with natural numbers $l$ and $k$, which is derived by restricting the optimization region of the measurement protocols. This bound $U_{l,k}$ is the generalization of the upper bound $U_{l}$, and is utilized in the numerical calculation in Sec. III.3.3. The definition and derivation of this type of bound are provided in Appendix D.
@@
where $\mathcal{N}_{n:l}$ denotes the set of measurement protocols composed of an arbitrary POVM (positive operator valued measurement) on the parties $\{a_{i}\}_{i=1}^{n}$ and local projective measurements on $l$ parties $\{a_{i}\}_{i=n+1}^{n+l}$  [^3] For $n$ such that $N-l+1\leq n<N$, $\mathcal{N}_{n:l}$ becomes the set of protocols where a POVM on $\{a_{i}\}_{i=1}^{n}$ and local projective measurements on $\{a_{i}\}_{i=n+1}^{N}$ are performed., and $Z_{n}$ is the outcome of the POVM, as shown in Fig. 6(c). By denoting the set of the local projective measurements on $\{a_{i}\}_{i=1}^{n+l}$ as $\mathcal{M}_{n+l}$, this bound is derived as follows:
@@
From Eq. (19), we can see that this lower bound $L_{l}$ is defined as the sum of the local terms $\min_{\Lambda}H_{\Lambda}(Y_{n:l}|Z_{n})$, which is similar to the upper bound $U_{l}$. The difference is that in the lower bound $L_{l}$, we need to optimize the POVM on the $d^{n}$-dimensional subsystem $\{a_{i}\}_{i=1}^{n}$, whose numerical cost seems to grow exponentially with the subsystem size $n$. However, in MPS with fixed bond dimension $D_{B}$, we can perform this optimization at a constant numerical cost, since the subsystem $\{a_{i}\}_{i=1}^{n}$ can be effectively regarded as a fixed-dimensional system ($D_{B}$-dimensional for open boundary conditions, $D_{B}^{2}$-dimensional for periodic boundary conditions). Because of this reduction of subsystem dimension, the numerical cost for $L_{l}$ becomes $Ne^{\mathcal{O}(l)}$ up to constant multiple in $N$-site MPS, and $e^{\mathcal{O}(l)}$ for the translationally invariant normal MPS in the thermodynamic limit.
@@
where $k$ is arbitrary natural number. Furthermore, for systems described with MPS, we can derive Eq. (14), which indicates that an arbitrary error bound is achieved with our method by increasing $l$. Since the numerical costs for calculating the bounds $L_{l}$ and $U_{l}$ scale as $e^{\mathcal{O}(l)}$ with the locality parameter $l$, Eq. (14) demonstrates that the cost scales with the inverse of the error bound, $\varepsilon$, as $e^{\mathcal{O}({\frac{1}{\varepsilon}})}$ (Table 1).
@@ III.3 Demonstration in examples
We demonstrate our efficient calculation method of work deficit density introduced in Secs. III.1 and III.2 in some examples. Especially, we address translationally invariant MPS in the thermodynamic limit, where the direct calculation is impossible, while our calculation method can be performed with reasonable numerical cost.
@@ III.3.1 AKLT state
We first calculate the work deficit density for the AKLT state Affleck et al. (1987, 1988) under open boundary condition. The AKLT state is the ground state of the AKLT Hamiltonian
@@
$A^{\uparrow}=\begin{pmatrix}0&0\\ -1&0\end{pmatrix},A^{0}=\sqrt{\frac{1}{2}}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},A^{\downarrow}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$ (22)
While the direct calculations of $\delta$ is infeasible since we need to optimize infinite sets of bases in the thermodynamic limit, the calculation of the lower bound $L_{1}$ can be performed with the optimization in $2\times 3$ dimensional system, due to the fixed bond dimension $D_{B}=2$. In this case, we can evaluate the lower bound through analytical calculation, as
@@ III.3.2 Cluster state
We next calculate the work deficit density in the cluster state Briegel and Raussendorf (2001) under open boundary condition. The cluster state is the ground state of the cluster Hamiltonian
@@
$A^{0}=\begin{pmatrix}0&0\\ 1&1\end{pmatrix},A^{1}=\begin{pmatrix}1&-1\\ 0&0\end{pmatrix}.$ (26)
While the direct calculation is not applicable in the thermodynamic limit, we can calculate the lower bounds $L_{2}$ by performing the optimization over $2\times 2\times 2$ dimensional system. This optimization can be performed analytically in this case, and the lower bound is obtained as
@@
Since the lower bound for $l=1$ is calculated as $L_{1}=0$ only providing the trivial bound, we can see that the bound is significantly improved by increasing locality parameter $l$.
We can also derive the upper bound $U_{\Lambda^{*}}$ with the ansatz protocol $\Lambda^{*}$ where the projective measurement on {|$\{\ket{0},\ket{1}\}$ and {$\{\ket{+},\ket{-}\}$ are alternately performed on each site. Here, basis $\ket{\pm}$ are defined as |±$\ket{\pm}\equiv\frac{1}{\sqrt{2}}(\ket{0}\pm\ket{1})$. This upper bound is calculated as
@@ III.3.3 MPS family under quantum phase transition
Figure 7: Correlation length as a function of the system parameter $g$ for the MPS family (Eq. (30)). The correlation length diverges at the quantum phase transition point $g=0$, indicating a transition from the SPT phase in $g<0$ to the trivial phase in $g>0$.
@@
We consider the region $-1\leq g\leq 1$, and system under preodic boundary condition. These MPS are the ground states of the parametrized Hamiltonian
@@
From Fig. 8(a), we can observe the bimodal structure of $\delta$, with its peaks in the regions $-1\leq g\leq 0$ and $0\leq g\leq 1$, respectively. The reasons for this bimodal structure can be attributed to two factors: long-range correlation around $g=0$ and the short-range correlation around $g=\pm 1$. At $g=1$, work deficit density $\delta$ becomes zero, since the system is in a product state and does not have inter-site correlation at all. On the other hand, at $g=0$, $\delta$ drop to zero due to the non-decaying correlation. The generalized GHZ state realized at $g=0$, is the superposition of the completely aligned states |$\ket{00...0}$ and |$\ket{11...1}$, and therefore the multipartite work deficit remains the same value $\Delta=\ln 2$ independent of the system size $N$. In the thermodynamic limit $N\to\infty$, this leads to $\delta\equiv\frac{\Delta}{N}\to 0$. Consequently, the peak of $\delta$ occurs between these two extremes, where the system has the intermediate-scale correlation length. The peak in the region $-1\leq g\leq 0$ occurs in a similar mechanism.
We further discuss the kinks observed in Fig. 8(a). In this figure, the unsmooth change in $\delta$ is observed at the phase transition point $g=0$, reflecting the drastic change of the quantum correlation structure. Around this point $g=0$, the optimal measurement protocol $\Lambda^{*}$ is kept unchanged; it is to perform the measurement on the basis {|$\{\ket{0},\ket{1}\}$ for all sites. On the other hand, the kink observed around $g=0.175$ in Fig. 8(a) is due to a discontinuous change in the optimal measurement protocol at this point; in the region $g<0.175$, the optimal protocol is to perform projective measurements on the basis {|$\{\ket{0},\ket{1}\}$ for all sites, while for $g\geq 0.175$, the optimal basis suddenly switches to {$\{\ket{+},\ket{-}\}$, where |±$\ket{\pm}\equiv\frac{1}{\sqrt{2}}(\ket{0}\pm\ket{1})$. Additionally, the kink may occur for the same reason in the region $\-1\leq g\leq 0$, although it is not clearly observable in the current numerical calculation due to the gap in the lower and upper bounds. Such kinks, resulting from changes in the optimal measurement protocols, do not necessarily indicate a drastic change in the quantum correlation structure at those points. To distinguish between such kinks and the unsmoothness due to the phase transition at $g=0$, we further introduce the framework of coarse-graining in Sec. IV, demonstrating that the kink around $g=0$ becomes sharper while the kinks at the other points shift or get dull under the coarse-graining.
@@ IV Coarse-graining
Figure 9: (a) Schematic for the coarse-graining. The neighbouring $m$ sites $\{a_{mn+1},a_{mn+2},\dots,a_{(m+1)n-1}\}$ are merged and regarded as a single party $b_{n+1}$. (b) Convergence of normal MPS to bipartite-entangled form under coarse-graining. By taking the party size $m$ much larger than the correlation length $\xi$, normal MPS converges to the form where bipartite-entangled pair $\sum_{i}\sqrt{q_{i}}\ket{i}\ket{i}$ is shared between the neighbouring parties.
In the foregoing sections, we have focused on the situation where each site is regarded as a single party. In this section, we calculate the work deficit density defined in a coarse-grained picture. Specifically, we consider the situation where $m$ sites $\{a_{mn+1},a_{mn+2},\dots,a_{(m+1)n-1}\}$ are merged and regarded as a single party $b_{n+1}$, as shown in Fig. 9(a). To represent the work deficit density defined under such a partitioning, we introduce the notation $\delta_{m}$, where the subscript $m$ denotes the number of sites per a single party. In the case of $m=1$, the subscript may be omitted, as has been done so far.
In the framework of coarse-graining, we can get access to the structure of longer-range multipartite quantum correlation. For example, the work deficit density $\delta_{m}$, defined under the partitioning of Fig. 9(a), does not account for the short-range quantum correlation among $m$ sites $\{a_{mn+1},\dots,a_{(m+1)n-1}\}$, while it captures the structure of longer-range quantum correlation well. This is contrary to $\delta_{1}$, which accounts for the effects of local quantum correlations across all sites. In this section, we analyze the long-range structure using the framework of coarse-graining, demonstrating that the quantum phase transition in the MPS family (30) can be effectively probed with this framework.